OpenStudy (anonymous):

What are the factors of x2 - 16 ?

OpenStudy (anonymous):

Recognize this is the difference of two squares: \[a^2 - b^2 = (a+b)(a-b)\] So therefore: \[x^2 - 4^2 = ?\]

OpenStudy (anonymous):

(x+4)(x-4)

OpenStudy (anonymous):

(x-4)(x+4)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

yeah